CVP with Activity-Based Costing and Multiple Products

Busy-Bee Baking Company produces a variety of breads. The plant manager would like to expand production into sweet rolls as well. The average price of a loaf of bread is $1. Anticipated price for a package of sweet rolls is $1.50. Costs for the new level of production are as follows:


Cost Driver
Unit Variable
Cost
Level of Cost
Driver
Loaf of bread $0.65           —          
Package of sweet rolls $0.93           —          
Setups $300           250          
Maintenance hours $15           3,500          
Other data:
Total fixed costs (traditional) $185,000  
Total fixed costs (ABC) 57,500  

Busy-Bee believes it can sell 600,000 loaves of bread and 200,000 packages of sweet rolls in the coming year.

Required:

1. Prepare a contribution-margin-based income statement for next year. Be sure to show sales and variable costs by product and in total.

Busy-Bee Baking Company
Contribution-Margin-Based Income Statement
Bread Sweet Rolls Total
Sales $ $ $
Less: Variable cost
Contribution margin $ $ $
Less: Fixed costs
Operating income $

Feedback

Remember a Contribution margin income statement calculates contribution-margin not gross profit.

2. Compute the break-even sales for the company as a whole using conventional analysis. Round your answer to the nearest dollar.
$

3. Compute the break-even sales for the company as a whole using activity-based analysis. Round your answer to the nearest dollar.
$

4. Compute the break-even units of each product in units. In your computations, round amounts to the nearest cent. Round your final answers to the nearest whole number of units.

Break-even loaves of bread
Break-even packages of sweet rolls

Does it matter whether you use conventional analysis or activity-based analysis?
Yes

5. Suppose that Busy-Bee could reduce the setup cost by $100 per setup and could reduce the number of maintenance hours needed to 1,000. How many units of each product must be sold to break even in this case? Round your answer up to the next higher whole unit (for example, 50.3 units rounds to 51). In your computations, round amounts to three decimal places.

Break-even loaves of bread
Break-even packages of sweet rolls
Answer #1
Part-1
Contributon Margin Based Income Statement
Bread Sweet Rolls Total
Unit Sold            6,00,000                       2,00,000
Sales @$1, $1.50            6,00,000                       3,00,000        9,00,000
Less: Variable Cost @$0.65, $0.93            3,90,000                       1,86,000        5,76,000
Contribution Margin            2,10,000                       1,14,000        3,24,000
Less: Fixed Cost        1,85,000
Net Income        1,39,000
Part-2
BEP Sales = Fixed Cost/ Contribution Margin Ratio
=185000/36%= $513889
CM Ratio= (324000/900000)=36%
Part-3
BEP Sales As per ABC
(( Fixed Cost + ( Setup Cost * No. of Setup)+( Maint. Cost X No. of Hour))/ CM Ratio
((57500+($300X 250)+($15X 3500))/ 36%= $513889
Part-4
Sales Mix Ratio of Bread ( 600000/9000000)=66.67%
Sales Mix Ratio of Sweet Roll ( 300000/9000000)=33.33%
BEP ( Unit )
Bread = Total BEP Sales X Sales Mix Ratio
$513889X 66.67%=$342610
Unit sells Price : $1
No. of BEP Unit= $342610/$1= 342610 unit
BEP ( Unit )
Bread = Total BEP Sales X Sales Mix Ratio
$513889X 33.33%=$171296
Unit sells Price : $1.50
No. of BEP Unit= $171296/$1.50= 114197 unit
Part-5
BEP Sales As per ABC
(( Fixed Cost + ( Setup Cost * No. of Setup)+( Maint. Cost X No. of Hour))/ CM Ratio
((57500+($200X 250)+($15X 1000))/ 36%= $3,40,278
BEP ( Unit )
Bread = Total BEP Sales X Sales Mix Ratio
$340278X 66.67%=$226863
Unit sells Price : $1
No. of BEP Unit= $226863/$1= 226863 unit
BEP ( Unit )
Bread = Total BEP Sales X Sales Mix Ratio
$340278X 33.33%=$113426
Unit sells Price : $1.50
No. of BEP Unit= $113426/$1.50= 75617 unit